OpenStudy (anonymous):

A ball is fired from the ground with an initial speed of 2.04 ✕ 103 m/s (approximately six times the speed of sound) at an initial angle of 63.0° to the horizontal. Neglect air resistance. (a) Find the ball's horizontal range. (b) Find the amount of time the ball is in motion.

OpenStudy (owlet):

projectile range is \(\LARGE \frac{v_i^2\ sin\ 2 \theta}{g}\) time of flight is \(\LARGE \frac{2v_i\ sin\ 2 \theta}{g}\) where \(v_i\) is the initial speed, theta is the given angle when the ball is launched and g is the acceleration due to gravity.